# Underage Detection through a Multi-Task and MultiAge Approach for Screening Minors in Unconstrained Imagery

### v1

@@ 3.2.1 Restricted Overall Test
The combined test bed of Section 3.1.1 includes the test sets of all age datasets used in this study (AFAD, AgeDB, CACD200, CLAP2016, Morph Album 2, UTKFace, FG-NET, the Dartmouth database of children’s faces, CASIA-Face-Africa). As described in Section 3.1.2, the data has been cleaned from label noise to some extent. We use the test split of Paplhám and Franc [15] (fold 0), except for FG-Net, Dartmouth, and CASIA-Face-Africa, which are not used in the training and are used entirely for testing. The overall test consists of 110k images ($30\%$ from AFAD and $34\%$ from CASIA-Face-Africa).
ASORES-39k: Age-estimation overall restricted test
The overall test is dominated by AFAD and CASIA-Face-Africa. This is problematic because both present challenging conditions. AFAD has tightly cropped faces of low image resolution, and its age labels seem to be quite noisy. At the same time, CASIA-Face-Africa has challenging imaging conditions of a low-quality webcam and an IR camera. Therefore, we define a Restricted Overall Test that excludes these two datasets, which will be considered for the ”wild” test case below. The restricted overall test consists of $39k$ images (38924), distributed across datasets as follows: AgeDB: $9\%$, CACD2000: $27\%$, CLAP2016: $5\%$, Dartmouth: $16\%$, FG-NET: $3\%$, MORPH: $28\%$, UTKFace: $12\%$.
3.2.2 ASWIFT-20k: Age-estimation wild-shifts test
Apart from overall high age regression and underage-detection accuracy, it is vital that the accuracy does not break down under real-world conditions, i.e., if variations in imaging conditions, poses, or facial expressions come into play. Therefore, we proposed ASWIFT-20k, a wild test, where we deliberately select images that deviate from standard conditions.
We inspected the $4.8k$ images of the UTKFace test set and manually selected $210$ images that showed interesting poses, facial expressions, and imaging conditions. Since manual selection is tedious and subjective, we turn to automatic selection criteria in the following criteria: UTK hard, Pose-45, expression, low-sharpness, low-contrast, low-brightness, low-saturation, and high-saturation.
Table 1 summarizes the composition of the ASWIFT-20k wild test in terms of source datasets according to the respective selection criteria.
@@ 3.2.2 ASWIFT-20k: Age-estimation wild-shifts test
Figure 3 shows the age distributions of the overall test and the restricted overall test (Section 3.2.1) and the wild test and its subsets (current section).
Figure 3: Age distribution (cumulative histogram) of the overall test, the restricted overall test, and the “wild” test and its subsets. Note that the “Pose-45” subset, i.e., those faces with pose angle beyond 45° come mostly from the Dartmouth Dataset of Children’s Faces and thus contains more children than adults.
@@ 3.2.3 Selection criteria implemented
Furthermore, we separate our test cases according to the imaging conditions. For all images of the overall test set, we have computed measures for brightness (mean of the grayscaled face patch), contrast (standard deviation of the grayscaled face patch), saturation (mean of the saturation channel in HSV color space), sharpness (variance of Laplacian) of a tight face crop.
For ASWIFT-20k, we select from the restricted overall test the lowest $8\%$ in contrast and sharpness, the lowest $4\%$ in brightness and saturation, and the highest $4\%$ in saturation (images with unnaturally high saturation may be out-of-distribution with respect to the training set, imposing difficulties for an age algorithm). We have added extreme imaging conditions from the AFAD test set and CASIA-Face-Africa data (only the webcam images, not the IR ones), set separately and with half the percentages mentioned above (otherwise, these would dominate the wild test too much). In total, $14683$ images were selected based on the imaging conditions (note that there is some overlap between the different criteria, e.g., many images are both low-contrast and low-sharpness or both low-contrast and low-brightness). Table 1 shows the breakdown into source datasets and selection criteria.
To work with the last criterion, facial expression, we use the MorphSet facial-expression estimator [^3] https://github.com/bbonik/facial-expression-analysis by Vonikakis et al. [28, 29], to single out extreme facial expressions from the test data and to define several respective test cases. Figure 4 gives a visual representation of the distribution of the selected face images in the arousal-valence space, together with the number of images selected for each expression category.
@@
The MorphSet model operates on frontalized landmarks obtained with the dlib library and provides expression estimates in the valence-arousal space via a partial least-squares regression (PLS). The valence axis ranges from negative/unpleasant to positive/pleasant, and the arousal axis from calm/passive to aroused/active, with neutral in the center. The Euclidean distance from the center is the intensity of the expression, and it is used as a criterion to select extreme expressions. We select all images with an intensity value above $0.7$ for most expressions in the “wild” test. For the expressions happy, pleased, and delighted, which are over-represented in the data, we apply higher thresholds. For the less common expressions gloomy, depressed, bored, droopy, tired, sleepy, calm, serene, content, and satisfied (all with negative arousal), we lower the thresholds such that 40 images are selected from the overall test.
@@ 4 Underage detection architecture with MultiAge network
In this work, we pick up the thread of the “FaRL + MLP” model introduced by Paplhám and Franc [15] and adapt it to the downstream task of non-adult detection under real-world conditions.
@@ 4.1 MultiAge network
Figure 5: Scheme of the model architecture: The face patch extracted from the image scene divided into subpatches (14$\times$14 non-overlapping patches of 16$\times$16 pixels), transformed with the FaRL image encoder [16] and then processed by our “MultiAge” network. Apart from an age estimate (with the typical estimation error), several outputs indicate whether the subject will likely be under a given age threshold. As discussed in Section 5 below, these underage outputs are more reliable than using the age estimate directly.
Figure 5 depicts a graphical summary of the underage detection architecture proposed. As illustrated in Fig. 5, an input face patch from an image scene is tokenized into patches of size 16x16, linearly projected and fed to the frozen FaRL image encoder [16], a ViT-B/16 transformer with $L{=}12$ blocks and hidden width $768$ [18]. The normalized and projected output of the FaRL image encoder is denoted $\mathbf{z}\!\in\!\mathbb{R}^{d}$, $d=512$ and constitutes the visual representation passed to the MultiAge network, a light MLP with $k=5$ task heads, one regressor and four binary detectors, which will be the only trainable part.
The code proposed by Paplhám and Franc [15], however, shows a deviation from this intended behavior: The dense layers are identical and share the same weights [^4] Details: https://github.com/paplhjak/Facial-Age-Estimation-Benchmark/issues/27, which limits the expressivity of the network and has a regularizing effect.
We investigate this weight–sharing found in Paplhám and Franc [15], expressed as Weight–sharing MLP (WS) against an independent two-layer design with different width configurations, we denoted as Independent MLP (IND). Let $m$ be the optional bottleneck width ($m\!\in\!\{128,256,512\}$). In the work of Paplhám and Franc [15], we could state that they proposed a Weight–sharing MLP as follows:
$\mathbf{h}_{\mathrm{WS}}=ReLU\!\bigl{(}\mathbf{W}\mathbf{z}+\mathbf{b}_{1}\bigr{)},$ (1)
$\mathbf{o}_{\mathrm{WS}}=ReLU(\mathbf{W}\mathbf{h}_{\mathrm{WS}}+\mathbf{b}_{2}),$
with a single $\mathbf{W}\!\in\!\mathbb{R}^{d\times d}$ reused twice. This reduces the parameter count from $2d^{2}$ to $d^{2}$ and regularizes the mapping, but constrains its rank to $d$.
@@
$\mathbf{h}_{\mathrm{IND}}=ReLU\!\bigl{(}\mathbf{W}_{1}\mathbf{z}+\mathbf{b}_{1}\bigr{)},\mathbf{W}_{1}\!\in\!\mathbb{R}^{m\times d},$ (2)
$\mathbf{o}_{\mathrm{IND}}=ReLU(\mathbf{W}_{2}\mathbf{h}_{\mathrm{IND}}+\mathbf{b}_{2}),\mathbf{W}_{2}\!\in\!\mathbb{R}^{d\times m}$
@@
In the baseline model [15], $\mathbf{o}$ is fed into to the age estimation head, implemented as classification in $c=102$ age classes for ages 0 to 101, trained with cross-entropy loss [15], and the predicted age $\hat{y}$ is given as
@@
To this point, the model estimates the age $\hat{y}$, and the age estimate can be used for non-adult detection, limited by the estimation errors.
In the spirit of multitask learning [30], we combine age estimation task with binary discrimination tasks of underage detection with the binary discriminators for over/under-age discrimination. On the one hand, these underage discriminators are of immediate use in applications (Fig. 5), on the other hand, this leverages synergies between the age-estimation task and the non-adult detection tasks, and the binary discrimination should help to guide the focus of the model to the age range of interest.
Thus, both for both cases of MLP architectures, $\mathbf{o}$ of Eq. 1 or Eq. 2 is forwarded to the age estimation head (Eq. 3) and to $k{-}1$ sigmoid-activated binary heads with thresholds $A=\{12,15,18,21\}$. The binary heads are given as
$u_{k}=\mathbf{w}_{k}^{\top}\mathbf{o}+b_{k},$ $\mathbf{w}_{k}\!\in\!\mathbb{R}^{d},$ (4)
$p_{k}=\sigma(u_{k}),$ $k=1,2,3,4,$
where $u_{j}$ and $p_{k}$ are the intermediate activation and final output of each head, respectively, and $\sigma(\cdot)$ is the sigmoid.
@@ 4.2 $\alpha$-reweighted focal loss and age-balanced resampling
Minors are underrepresented in the training set, which tends to bias the trained model towards predicting that subjects are adults. This is detrimental to our objective of faithfully detecting non-adults (and faithfully estimating the age of minors). To mitigate this imbalance, we propose a variation of the focal loss proposed by Lin et al. [31], an $\alpha$-reweighted focal loss variant to prioritize minority class examples.
In case there are several non-adult detectors with different age levels, these have different class imbalances and different values of the correcting $\alpha$ parameter.
@@
$\mathcal{L}_{\alpha}(y,\hat{y})=-\alpha_{y}(1-\hat{y})^{\gamma}y\log(\hat{y})-(1-\alpha_{y})\hat{y}^{\gamma}(1-y)\log(1-\hat{y}),$ (5)
where $\hat{y}$ is the predicted probability, $y$ is the binary ground-truth, and $\gamma$ is the focusing parameter (set to $\gamma=2$ following [31], for $\alpha=1$, $\gamma=0$, Eq. 5 reduces to the binary cross-entropy loss). The weighting term $\alpha_{y}$ we propose is computed as follows:
$\alpha_{y}=\begin{cases}\sqrt{\frac{N_{\text{majority}}}{N_{\text{minority}}}},&\text{if }y=1\text{ (minority)},\\[6.0pt] \sqrt{\frac{N_{\text{minority}}}{N_{\text{majority}}}},&\text{otherwise.}\end{cases}$ (6)
@@
An et al. [32] have shown that, for training methods like stochastic gradient descent, reweighting the training losses can have detrimental effects. Thus, we additionally proposed an age-balanced mini-batch resampling scheme. We employ an age-equalized training scheme in which the age label determines the statistical weight for each image to be selected for the current minibatch during training. The age range is partitioned into $n$ age groups, concretely $n=12$, with age boundaries at the ages 4, 8, 12, 16, 20, 24, 28, 32, 36, 42, and 50. Then, the weight factors for each age group are chosen inversely proportional to the number of images in the age group.  [^5] Resampling changes the meaning of the training epoch, which means that as many training patterns have been drawn as the size of the training set, but some will not have occurred in this epoch and more than once. Formally, the probability of sampling from age-bin $i$ is set as:
@@ 4.3 Mitigating the impact of edge cases
Considering the binary discrimination task of non-adult detection for a given age threshold as a function of age, it is clear that misclassifications are most likely to happen for subjects close to the age threshold. For example, the age estimate for a subject of $17$ years and $11$ months may be easily confused with that of a subject of $18$ years and one month. At the same time, these edge cases may produce strong gradients and confuse the model during training. Therefore, we propose an age-gap hyperparameter, defining an interval around the age threshold; patterns with age in the age gap are neither positive nor negative and do not contribute to the loss of the respective discriminator.
@@
where $a$ denotes the actual age. The size of the gap $G$ is a parameter to be determined empirically through experiments (below, In Section 5.4.2 we will observe optimal balance between discriminative power and gradient stability at at $G\approx T/6$). On the one hand, larger age gaps excessively remove informative training examples. On the other hand, smaller gaps may provide insufficient protection against noisy gradients during training. We hypothesize that this approach may effectively reduce training ambiguity, improving the optimization stability, and generalization around critical legal thresholds.

### v2

@@ 3.2.1 Restricted Overall Test
The combined test bed of Section 3.1 includes the test sets of all age datasets used in this study (AFAD, AgeDB, CACD200, CLAP2016, MORPH Album 2, UTKFace, FG-NET, the Dartmouth database of children’s faces, CASIA-Face-Africa). As described at the end of Section 3.1, the data has been cleaned from the worst label noise. We use the test split from Paplham_CVPR2024_benchmark (fold 0), except for FG-NET, Dartmouth, and CASIA-Face-Africa, which are not used in the training and are used entirely for testing. The overall test consists of 110k images ($30\%$ from AFAD and $34\%$ from CASIA-Face-Africa).
ASORES-39k, the “Age Estimation Overall Restricted Test”
AFAD and CASIA-Face-Africa dominate the overall test. This is problematic because both present challenging conditions. AFAD has faces cropped tightly and at low resolution, and its age labels seem noisy. At the same time, CASIA-Face-Africa has challenging imaging conditions, including a low-quality webcam and an IR camera. Therefore, we define a Restricted Overall Test that excludes these two datasets, which will be considered for the ”wild” test case below. The restricted overall test consists of $39k$ images (38924), distributed across datasets as follows: AgeDB: $9\%$, CACD2000: $27\%$, CLAP2016: $5\%$, Dartmouth: $16\%$, FG-NET: $3\%$, MORPH: $28\%$, UTKFace: $12\%$.
3.2.2 ASWIFT-20k, the unconstrained “Age Estimation Wild-Shifts Test”
Apart from overall high age regression and underage detection performance, it is vital that the accuracy does not break down under real-world conditions, i.e., if variations in imaging conditions, poses, or facial expressions come into play. Therefore, we propose ASWIFT-20k, a wild test, where we deliberately select images that deviate from standard conditions. We inspected the $4.8k$ images of the UTKFace test set and manually selected $210$ images that showed interesting poses, facial expressions, and imaging conditions. Since manual selection is tedious and subjective, we turn to automatic selection criteria by pose angles facial expression and image properties. The selection criteria are described in detail in the following subsection 3.2.3.
@@ 3.2.2 ASWIFT-20k, the unconstrained “Age Estimation Wild-Shifts Test”
The resulting composition of the ASWIFT-20k wild test in terms of source datasets and selection criteria is summarized in Table 1. Figure 2(c) shows the age distributions of the overall test, the ASORES-39k test and the ASWIFT-20k test and its subsets. Finally, Figure 3 shows the demographic distributions of the composed test sets ASORES-39k and ASWIFT-20k as heat maps. While ASWIFT-20k has is wider distribution in the ethnicity dimension than ASORES-39k, both show strong signatures of their respective constituents: blacks are males aged 18 to 60 years (from MORPH), white children of ages 6 to 9 (from Dartmouth), and Asians aged 18 to 30 years (from AFAD). These imbalances have to be kept in mind when analyzing performance measures on these test sets.
Figure 3: Distribution of ASORES-39k and ASWIFT-20k over gender, ethnicity (FairFace [Karkkainen_2021_FairFace]) and age.
@@ 3.2.3 Selection criteria implemented
Furthermore, we separate our test cases by imaging conditions. For all images of the overall test set, we have computed measures for brightness, contrast (mean and standard deviation of the grayscaled face patch, respectively), saturation (mean saturation in HSV color space) and sharpness (variance of Laplacian) of a tight face crop. For ASWIFT-20k, we select from the restricted overall test the lowest $8\%$ in contrast and sharpness, the lowest $4\%$ in brightness and saturation, and the highest $4\%$ in saturation. We have added extreme imaging conditions from the AFAD test set and the CASIA-Face-Africa dataset (only webcam images, not the IR ones) separately and with half the percentages mentioned above (otherwise, these would dominate the wild test too much). In total, $14683$ images were selected based on the imaging conditions (note that there is some overlap across the different criteria; e.g., many images are both low-contrast and low-sharpness, or both low-contrast and low-brightness). Table 1 shows the breakdown into source datasets and selection criteria.
To work with the last criterion, facial expression, we use the MorphSet facial-expression estimator by Vonikakis et al. [Vonikakis_2021_expression], to single out extreme facial expressions. The MorphSet model operates on frontalized landmarks obtained with the dlib library and provides expression estimates in the valence-arousal space via a partial least-squares regression. The valence axis ranges from negative/unpleasant to positive/pleasant, and the arousal axis from calm/passive to aroused/active, with neutral in the center. Figure 4 gives a visual representation of the distribution of the selected face images in the arousal-valence space, together with the number of images chosen for each expression category.
@@
The Euclidean distance from the center is the intensity of the expression, and it is used as a criterion to select extreme expressions. We select all images with an intensity value above $0.7$ for most expressions in the “wild” test. For the expressions happy, pleased, and delighted, which are over-represented in the data, we apply higher thresholds. For the less common expressions gloomy, depressed, bored, droopy, tired, sleepy, calm, serene, content, and satisfied (all with negative arousal), we lower the thresholds such that 40 images are selected.
@@ 4 Underage detection architecture with MultiAge network
In this work, we pick up the thread of the “FaRL + MLP” model introduced by Paplham_CVPR2024_benchmark and adapt it to the downstream task of underage detection under real-world conditions. With 87M parameters, the image encoder of FaRL [Zheng_2022_FaRL], a 12-layer vision transformer [Dosovitskiy_2021_ViT], is large compared to early CNNs and “slim” models like SSR-Net [Yang_2018_SSRNet], but still smaller than the VGG-16 of Rothe_2018_IMDBWIKI.
@@ 4.1 MultiAge network
Figure 5: Scheme of the model architecture: The face patch extracted from the image scene is divided into subpatches ($14\times 14$ non-overlapping patches of $16\times 16$ pixels), transformed with the FaRL image encoder [Zheng_2022_FaRL] and then processed by our “MultiAge” network. Apart from an age estimate (with its typical estimation error), several outputs indicate whether the subject is likely to be under a given age threshold. As discussed in Section 5 below, these underage outputs are more reliable than using the age estimate directly.
Figure 5 depicts a graphical summary of the underage detection architecture proposed. An aligned input face patch ($224\times 224$ pixels) is extracted from an image scene. During training, this is done with random augmentations, randomized crop, and horizontal flip. The face patch is tokenized into patches of size $16\times 16$, linearly projected and fed to the frozen FaRL image encoder [Zheng_2022_FaRL], a ViT-B/16 transformer with L= 12L\,{=}\,12 blocks and hidden width $768$ [Dosovitskiy_2021_ViT]. The normalized and projected output of the FaRL image encoder is denoted $\mathbf{z}\!\in\!\mathbb{R}^{d}$, $d=512$, and constitutes the visual representation passed to the MultiAge network, a light MLP with $k=5$ task heads, one regressor and four binary detectors, which will be the only trainable part.
The code proposed by Paplham_CVPR2024_benchmark, however, shows a deviation from this intended behavior: the dense layers are identical and share the same weights [^3] Details: https://github.com/paplhjak/Facial-Age-Estimation-Benchmark/issues/27, which limits the expressivity of the network and has a regularizing effect. We investigate this weight–sharing found in Paplham_CVPR2024_benchmark, expressed as Weight–sharing MLP (WS) against an independent two-layer design with different width configurations, we denoted as Independent MLP (IND). In the work of Paplham_CVPR2024_benchmark, they implemented a weight–sharing MLP as follows:
$\mathbf{h}_{\mathrm{WS}}$ $=ReLU\!\bigl(\mathbf{W}\,\mathbf{z}+\mathbf{b}_{1}\bigr),$ $\mathbf{o}_{\mathrm{WS}}$ $=ReLU(\mathbf{W}\,\mathbf{h}_{\mathrm{WS}}+\mathbf{b}_{2}),$ (1)
with a single $\mathbf{W}\!\in\!\mathbb{R}^{d\times d}$ used twice. This reduces the parameter count from $2d^{2}$ to $d^{2}$ and regularizes the mapping, but constrains its rank to $d$.
@@
$\mathbf{h}_{\mathrm{IND}}$ $=ReLU\!\bigl(\mathbf{W}_{1}\,\mathbf{z}+\mathbf{b}_{1}\bigr),\mathbf{W}_{1}\!\in\!\mathbb{R}^{m\times d},$ $\mathbf{o}_{\mathrm{IND}}$ $=ReLU(\mathbf{W}_{2}\,\mathbf{h}_{\mathrm{IND}}+\mathbf{b}_{2}),\mathbf{W}_{2}\!\in\!\mathbb{R}^{d\times m}$ (2)
@@
In the baseline model [Paplham_CVPR2024_benchmark], $\mathbf{o}$ is fed into the age estimation head, implemented as classification in $c=102$ age classes for ages 0 to 101, trained with cross-entropy loss [Paplham_CVPR2024_benchmark], and the predicted age $\hat{y}$ is given as
@@
To this point, the model estimates the age $\hat{y}$, which can be used for underage detection, though it is limited by estimation errors.
In the spirit of multitask learning [Caruana_1997_Multitask], we combine age estimation task with binary discrimination tasks of underage detection with the binary discriminators for over/underage discrimination. On the one hand, these underage discriminators are of immediate use in applications (Fig. 5), on the other hand, this leverages synergies between the age estimation task and the underage detection tasks, and the binary discrimination should help to guide the focus of the model to the age range of interest.
Thus, for both cases of MLP architectures, $\mathbf{o}$ of Eq. 1 or Eq. 2 is forwarded to the age estimation head (Eq. 3) and to k1k{-}1 sigmoid-activated binary heads with thresholds $A=\{12,15,18,21\}$. The binary heads are given as:
$u_{k}=\mathbf{w}_{k}^{\top}\mathbf{o}+b_{k},$ $\mathbf{w}_{k}\!\in\!\mathbb{R}^{d},\quad$ $p_{k}=\sigma(u_{k}),$ $k=1,2,3,4,$ (4)
where $u_{k}$ and $p_{k}$ are the intermediate activation and final output of each head, respectively, and $\sigma(\cdot)$ is the sigmoid.
@@ 4.2 $\alpha$-reweighted focal loss and age-balanced resampling
Minors are underrepresented in the training set, which tends to bias the trained model towards adults. This is detrimental to our objective of faithfully detecting minors and estimating their age. To mitigate this imbalance, we propose a variation of the focal loss of Lin et al. [Lin_ICCV_2017_FocalLoss], an $\alpha$-reweighted focal loss variant to prioritize minority class examples. In the case of several underage detectors with different age levels, they exhibit different class imbalances and different values of the correction parameter $\alpha$.
@@
[…]
where $\hat{y}$ is the predicted probability, $y$ is the binary ground-truth, and $\gamma$ is the focusing parameter (set to $\gamma=2$ following [Lin_ICCV_2017_FocalLoss], for $\alpha=1$, $\gamma=0$, Eq. 5 reduces to the binary cross-entropy loss). The weighting term $\alpha_{y}$ we propose is computed as follows:
$\alpha_{y}=\begin{cases}\sqrt{{N_{\text{majority}}}/{N_{\text{minority}}}},&\text{if }y=1\text{ (minority)},\\[6.0pt] \sqrt{{N_{\text{minority}}}/{N_{\text{majority}}}},&\text{otherwise.}\end{cases}$ (6)
@@
An_2021_Resampling have shown that, for training methods like stochastic gradient descent, reweighting the training losses can have detrimental effects. Thus, we additionally propose an age-balanced mini-batch resampling scheme. We employ an age-equalized training scheme in which the age label determines the statistical weight for each image to be selected for the current minibatch during training. The age range is partitioned into $n$ age groups, specifically $n=12$, with age boundaries at 4, 8, 12, 16, 20, 24, 28, 32, 36, 42, and 50. Then, the weight factors for each age group are chosen inversely proportional to the number of images in the age group. Formally, the probability of sampling from age-bin $i$ is set as:
@@ 4.3 Mitigating the impact of edge cases
Considering the binary discrimination task of underage detection for a given age threshold as a function of age, it is clear that misclassifications are most likely to happen for subjects close to the age threshold. For example, the age estimate for a subject of $17$ years and $11$ months may be easily confused with that of a subject of $18$ years and one month. At the same time, these edge cases may produce strong gradients and confuse the model during training. Therefore, we propose an age-gap hyperparameter, defining an interval around the age threshold; patterns with age in the age gap are neither positive nor negative and do not contribute to the loss of the respective discriminator. Note, however, that due to the multitask setup and the shared representation $\mathbf{o}$, examples from the age gap still contribute indirectly to the training of the respective underage detection task.
@@
where $a$ denotes the actual age. The value of $G$ is a hyperparameter that has to be validated empirically. Too large age gaps would remove too many informative training examples, while too small gaps may provide insufficient protection against noisy gradients during training.
The facial appearance changes more rapidly at young ages than at older ages [Geng_2014_ALDL]. Therefore, it is natural to define the age gap in relative terms with a scale parameter $f>1$, excluding a gap from $T/f$ to $Tf$ for the under-$T$ detector. Empirically, we find optimal results with $f=6/5$ (see Section 5.4.2 below). Taking into account the integer nature of the age labels through rounding towards zero, this results in the age gaps
$12+[-2,+2],\quad 15+[-2,+3],\quad 18+[-3,+3],\quad 21+[-3,+4]$ (9)
for $T=12,15,18,21$, respectively.